OpenStudy (anonymous):

WILL GIVE MEDAL! Write the standard form of: y^2+4y-8x+4=0

OpenStudy (marigirl):

Can you please give more information.. ?

OpenStudy (anonymous):

This is is a parabola that I need to put in standard form

OpenStudy (marigirl):

well usually a standard form of a parabola is y=ax2+bx+c right? so i am guessing its asking us to make y the subject? I am not too sure.... how about \[(y+2)^2-8x=0\]

OpenStudy (marigirl):

or

OpenStudy (marigirl):

i think the second one could be more appropriate for a standard form

OpenStudy (anonymous):

I got (y+2)^2=8x for the standard form, so its the same

OpenStudy (marigirl):

yes, but (y+2)^2=8x f that kind of looks like a vertex form too doesn't it?

OpenStudy (anonymous):

a bit, but that's what I got and a online website

OpenStudy (marigirl):

which shows the y intercept....so it could be regarded as a vertex form (turning point form) what do u think

OpenStudy (marigirl):

Keep both answers as a reference if possible.

OpenStudy (anonymous):

Ok, do you know anything about finding intersection points with a system of equations?